Amber Company produces iron table and chair sets. During October, Ambers costs were as follows: $2.20 per Ib. $7.40 per hour4. Compute the standard direct labor rate for October. (Round your answer to 2 decimal places.) Standard Direct Labor Rate 5.
Answer #1

1. Total purchases = 1105000*2.20 = $2431000

2. Direct material price variance = (2-2.20)*1105000 = $221000 Unfavorable

3. Material quantity variance = (960000-990000)*2 = $60000 Unfavorable

4. Direct labor price variance = (standard rate - actual rate)*actual hours

5400 = (Standard direct labor rate - 7.40)*9000

Standard direct labor rate = $8 per hour

5. Direct labor efficiency variance = (10000-9000)*8 = $8000 Favorable